OpenStudy (anonymous):

Find the slope of the line and sketch the graph 3x-3y=18

OpenStudy (vishal_kothari):

3x-3y=18 -3y=-3x+18 y=1x-6 y=mx+b m is slope b is y-intercept the slope is 1 because I subtract 3x by both sides and divide -3 both sides...